# Large Language Models Synergize with Automated Machine Learning

### v2

@@ 6.4 Improvements from optimization
Table 3: Mean rank of the identified ML program in 10 trials. (1) Traditional: tasks based on traditional ML algorithms such as random forest and SVM. (2)Rows with white and grey backgrounds represent the search with cost@25 and cost@100 respectively.
@@ 7.1 Synergy between LLMs and autoML
A major goal of autoML is to democratize ML by reducing the amount of effort required by human experts (Thornton et al., 2013; Wang et al., 2021a; Xin et al., 2021; Van der Blom et al., 2021). AutoML has achieved substantial progress in recent years on an algorithm level (Zoph & Le, 2016; Liu et al., 2018). However, in practice, autoML methods are gradually becoming a tool for experts to potentially boost the performance of the workflow (Tornede et al., 2023; Van der Blom et al., 2021), due to the following three issues: (1) most autoML methods only select the solutions at the algorithm level. However, the programming required for implementing the algorithms can demand substantial human effort and expertise. (2) for deep learning, many autoML methods select model architectures, such as the number of convolutional layers, from scratch and train the models anew, which is prohibitively expensive for regular users in practical scenarios under low resource constraints (Öztürk et al., 2022). (3) many autoML methods do not offer a comprehensive and unified solution applicable across task modalities (e.g., traditional algorithms, CV and NLP).
In our study, autoML numerically evaluates and selects ML programs generated by LLMs. Conversely, LLMs offer a solution to all three issues of autoML. As demonstrated in Figure 1, Section 6, and Figure 9, LLMs can generate the programs for the entire ML workflow across task modalities, while directly utilizing pretrained models for deep learning. Without the code generation from LLMs, there would be no code for autoML to select and optimize in the first place.
@@ 7.2 Sequence generation and autonomous agent
Our method, Text-to-ML, seamlessly integrates all intermediary steps into a single fully autonomous process (Figure 3 and Algorithm 2). Future research may explore such multi-step seq2seq tasks as autonomous agents. Another interesting future exploration is the incorporation of additional text-to-text or image-to-text methods to automatically generate textual task descriptions, further automating the process.
@@ 7.3 Limitations
While the complexity of Contextual Modular Generation could increase only linearly over the problem length (Theorem 3.8), implementing this framework demands considerable effort (Appendix C.1). Furthermore, our approach still struggles at highly challenging ML tasks, particularly those that involve highly complex forms of input and output data (Appendix D). Whether addressing these tasks requires a more efficient implementation of the framework or a transition to a different framework remains an open problem.
@@ 8 Conclusion
The testing technique Constrained Generative Unit Testing is instrumental for the implementation of the framework. As a distinctive advantage of Text-to-ML, the compatibility with pre-written modules, including a consistent numerical evaluation procedure, enables significant increases in the performance of the programs via numerical optimization from autoML. In addition, we demonstrate the effectiveness of ZC proxies for pretrained models in both CV and NLP.
@@ A.1 Large Language Models
Iterative methods repeatedly generate the output sequences in order to improve the output sequence based on a sequence evaluation procedure. Let the initial input sequence be $X_{0}$, the sequence generation process be $f_{\lambda}$, and the evaluation procedure be $\mathbb{E}$. Then the output sequence in the $i_{th}$ iteration is $Y_{i}=f_{\lambda}(\Phi_{i}(Y_{<i},\mathbb{E}(Y_{<i}),X_{0}))$, where $\Phi_{i}$ represents a mechanism that assembles the input sequence for the iteration by integrating the output sequences, their evaluations, and the initial input sequence from preceding iterations. Recent works in this line include Yao et al. (2022); Madaan et al. (2023); Cobbe et al. (2021); Shinn et al. (2023); Chen et al. (2023b). In a reinforcement learning context, each output sequence represents an action (Yao et al., 2022; Shinn et al., 2023). The implementation of our method incorporates the self-reflection mechanism in Shinn et al. (2023). More specifically, $\Phi_{i}$ in our implementation assembles input sequence from code execution results and an explanation of the results.
Prompting methods improves the output sequence $Y$ by transforming the input sequence $X$. Let $\psi$ represents the prompt engineering process, then $p(Y|X)=\prod_{n=1}^{N}p_{\lambda}(y_{n}|y_{<n},\psi(X))$. Notable prompting methods including in-context learning (Wei et al., 2022; Zhou et al., 2022a; Wang et al., 2022; Min et al., 2022; Yao et al., 2023; Li et al., 2023) and zero-shot prompting (Kojima et al., 2022). Notably, although some prompting methods break the problem into intermediate steps, such as Wei et al. (2022) and Li et al. (2023), these methods perform the steps in one sequence generation process, differing from our definition of compositional reasoning (see Section 2.1).
@@ A.2 Automated Machine learning
Table 4: Comparison with existing methods in terms of qualitative features. (1) Code generation: whether the method automatically generates programs (2) Deep learning: whether the method automatically generates (part of) programs for deep learning tasks. (3) Data preparation: whether the method automatically chooses the strategy for data preparation (4) Modeling: whether the method automatically chooses the strategy for modeling (5) Optimization: whether the method optimizes the programs (components). * Although the method can automatically choose the strategy for modeling, the method does not automatically generate the program for modeling. ** The method uses the LLM as a pseudo-optimization model instead of a dedicated optimization algorithm (e.g., evolutionary computation and reinforcement learning.)
@@ C.1 Unit Testing Details
In this context, distinguishing between the input/output in the ML task and for the modules is crucial. For instance, the input to the data preparation module consists of raw data, while its output comprises the processed input data (features) and output data (targets) in the ML task. The unit tests described in Algorithm 3 focus on the ML task’s input and output data by scrutinizing the output produced by each module when provided with its respective input. Apart from the data preparation module, all other modules are given synthetic input/output data in the ML task as input.
@@
In our experiments, we utilize the Python libraries Pytorch, PyTorch Lightning, and Transformers for deep learning tasks and Scikit-learn, XGBoost, CatBoost, and LightGBM for tasks involving tabular data. For deep learning, the specifications in the plan include (1) the count of tensor types for each of input and output in the ML task. (2) the dimensionality of each tensor type. (3) the meaning of each dimension. (4) the size of each dimension. (5) isomorphic dimensions: dimensions that should share the same size over different settings of the data preparation process and modeling configuration. (6) suitable ranges of the size for each dimension. For tabular data tasks, the specifications are limited to the dimensionality of arrays for each of input and output in the ML task.
@@ C.3 Search Strategy
In our study, we employ two search strategies: random search and BOHB (Falkner et al., 2018). Random search is a powerful baseline in hyperparameter optimization due to its wide applicability and search efficiency for tasks with low effective dimensionality (Bergstra & Bengio, 2012). BOHB combines Bayesian Optimization (BO) and Hyperband (HB) (Li et al., 2018) to optimize hyperparameters efficiently. BOHB inherits HB’s robustness and flexibility, and BO’s ability to ensure strong performance. BOHB starts by rapidly identifying promising configurations like HB and refines them using BO to achieve optimal solutions faster. It’s designed to be simple, computationally efficient, and adaptable to various tasks. For BOHB, we use the default hyperparameters in Falkner et al. (2018) and set 30 epochs as the maximum budget for deep learning tasks.
@@ C.4 Module Generation
In our study, for each ML task, we generate three modules: data preparation, modeling, and post processing. Every module is generated through a step (${j}$) in Algorithm 1. First, the choice specifications (e.g., choice of feature engineering method or choice of models) from Search Space Generation are given as input sequence $X^{j}$. Subsequently, the LLM $f_{\lambda}$ generates the module $Y^{j}$. Then the generated module $Y^{j}$ is assessed with the contextual modular evaluation $\mathbb{E}^{j}$, which is the unit test for the module generated from Algorithm 3. After the tests, Module Generation employs the self-reflection strategy from Shinn et al. (2023) to improve the generation efficiency. The input sequence assembler $\Phi^{j}$ instructs the LLM $f_{\lambda}$ to reflect on the feedback from the tests and then assemble previously generated codes, the tests feedback, and the reflection into the following form for the next iteration:
@@ C.5 ZC proxies details
naswot Mellor et al. (2021) leverages the overlap of activations between data points in untrained networks to derive a measure indicative of the network’s future trained performance. This approach allows for rapid and efficient searching of powerful networks without the extensive computational costs typically associated with training each candidate network during the search process. Originally designed for network pruning, synflow identifies sparse trainable subnetworks within untrained neural networks. The key principle of synflow is to maintain a balance of synaptic strengths through the network, thus preserving the network’s ability to be trained post-pruning.
Except for params, the ZC proxies require passing data to the models. To investigate the 3 ZC proxies’ applicability with synthetic data, we first generate the program for producing synthetic data, then sample 50 versions of synthetic data based on different random seeds with a fixed shape of the data. We find the variations in the performance metrics reported by each proxy are minimal. Notably, synflow exhibits the most significant variation, with a Mean Relative Deviation (MRD) of $0.76\%$, indicating the suitability of the ZC proxies for synthetic data.
@@ C.7 Hyperparameters for finetuning
Our research focuses on an unusual search space with pretrained models at the architectural level. Öztürk et al. (2022) also explored a similar search space. Consequently, we utilize a comparable set of hyperparameters for fine-tuning the models (Table 5). To tailor this approach to our specific requirements, we have reduced the minimum batch size from 16 to 2 to accommodate the varying GPU memory demands across different tasks.

### v3

@@ 6.2 Generation of valid programs
For the code generation process, we analyze the mean number of iterations required to correct the errors (Table 8 in Appendix F). In the experiments, for code that fails unit tests, we calculate the average number of iterations needed to pass the tests again. We only count corrections before the forced memory reset. Overall, (1) data preparation modules are generally the most difficult to correct, averaging 5.43 iterations; (2) for GPT-4, with its larger input token limit, the self-reflection mechanism shows the greatest benefit; (3) for LLMs with smaller input token limits (GPT-3.5 and PaLM 2), the memory from self-reflection can become unmanageable when a single previous iteration exceeds the input token limit, resulting in automatic memory truncation.
@@ 6.4 Improvements from optimization
Table 3: Mean rank of the identified ML program in 10 trials. (1) Traditional: tasks based on traditional ML algorithms such as random forest and SVM. (2)Rows with white and grey backgrounds represent search results with cost@25 and cost@100 respectively.
@@ 7.1 Synergy between LLMs and autoML
A major goal of autoML is to democratize ML by reducing the amount of effort required by human experts (Thornton et al., 2013; Wang et al., 2021a; Xin et al., 2021; Van der Blom et al., 2021). AutoML has achieved substantial progress in recent years on an algorithm level (Zoph & Le, 2016; Liu et al., 2018). However, in practice, autoML methods are gradually becoming a tool for experts to potentially boost the performance of the workflow (Tornede et al., 2023; Van der Blom et al., 2021), due to the following three issues: (1) most autoML methods only select the solutions at the algorithm level. However, the programming required to implement the algorithms can demand substantial human effort and expertise. (2) for deep learning, many autoML methods select model architectures, such as the number of convolutional layers, from scratch and train the models anew, which is prohibitively expensive for regular users in practical scenarios under low resource constraints (Öztürk et al., 2022). (3) many autoML methods do not offer a comprehensive and unified solution applicable across task modalities (e.g., traditional algorithms, CV and NLP).
In our study, autoML numerically evaluates and selects ML programs generated by LLMs. Conversely, LLMs offer a solution to all three issues of autoML. As demonstrated in Figures 1 and 9 and discussed in Section 6, LLMs can generate the programs for the entire ML workflow across task modalities, while directly utilizing pretrained models for deep learning. Without the code generation from LLMs, there would be no code for autoML to select and optimize in the first place. Our approach not only generates this code but also ensures that the code is natively compatible with optimization algorithms.
@@ 7.2 Sequence generation and autonomous agent
Our method, Text-to-ML, seamlessly integrates all intermediary steps into a single fully autonomous process (Figure 3 and Algorithm 2). Future research may explore such multi-step seq2seq tasks as autonomous agents. Another interesting future exploration involves incorporating text-to-text or image-to-text methods to enhance task description generation, further automating the process.
@@ 7.3 Limitations
While the complexity of Contextual Modular Generation could increase only linearly over the problem length (Theorem 3.8), implementing this framework demands considerable effort (Appendix C.1). The practical significance of the theoretical linear complexity should be more extensively tested for other types of sequence generation tasks. Furthermore, our approach still struggles with highly challenging ML tasks, particularly those that involve highly complex forms of input and output data (Appendix D). Whether addressing these tasks requires a more efficient implementation of the framework or a transition to a different framework remains an open problem.
@@
To target performance enhancement in a competitive environment, a future direction could involve significantly expanding the number of candidate solutions. For example, in our method, a modeling module might consist of a pretrained model and a custom architecture designated by the LLMs (Appendix C.2, Figure 1 and Figure 9). Future research could extend Text-to-ML with additional autoML techniques, such as predictor-based methods (Dudziak et al., 2020) and weight-sharing methods (Pham et al., 2018), to search for the custom architecture. Whether and to what extent such a pursuit can coexist with applications for regular users under low resource constraints is a crucial consideration. Additionally, establishing a systematic benchmark to compare automated solutions with human-generated ones is another significant challenge in this direction.
@@ 8 Conclusion
The testing technique Constrained Generative Unit Testing is instrumental for the implementation of the framework. A notable advantage of our approach is the seamless compatibility between generated and pre-written modules, allowing for consistent numerical evaluation and native integration with optimization algorithms. This compatibility enables autoML-driven numerical optimization to substantially improve program performance. In addition, we demonstrate the effectiveness of ZC proxies for pretrained models in both CV and NLP.
Broader Impact Statement
From the perspective of autoML, our approach aims to enhance the automation of the ML workflow, especially for everyday scenarios with limited resources. Text-to-ML is a potential tool for democratizing ML by significantly reducing the required efforts, expertise, and computational resources. Users simply provide a textual description of the task, and from there, the process is fully automatic, making ML in practice accessible to a broader audience.
However, the broader impacts of Text-to-ML must be carefully considered to ensure its responsible and ethical use. Our commitment to responsible AI is paramount, and we strongly urge the responsible use of automated tools for ML, including Text-to-ML.
Accessible Machine Learning: Benefits and Risks
By enabling a broader range of individuals and organizations to leverage ML, the Text-to-ML framework can foster innovation, accelerate scientific discovery, and enhance decision-making across various fields. However, this increased accessibility also introduces potential risks, such as misuse by individuals with malicious intent or those lacking a proper understanding of ML principles. We list the guidelines for responsible use of Text-to-ML as below:
  • Ethical Education: Users should complete a general ethics course in ML/AI before accessing Text-to-ML. This course would cover topics including fairness, bias, and risks in ML, and provide examples of both responsible and irresponsible uses of ML technology.
  • Bias and Toxicity in Data: Before applying Text-to-ML to a dataset, we recommend thoroughly checking for common biases such as selection bias, label bias, and sampling bias to ensure the dataset represents diverse populations, avoids subjective influences, and uses appropriate sampling methods (Mehrabi et al., 2021). Additionally, it is crucial to filter out toxicity in the data, including hate speech, offensive content, and misinformation.
  • Data Privacy: Users of Text-to-ML must adhere to relevant data privacy laws and regulations, such as the General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) in the European Union and similar laws in other jurisdictions. For instance, using Text-to-ML on personal data without consent or scraping data from unauthorized sources would be strictly prohibited.
  • Application Restrictions: The use of Text-to-ML for irresponsible or harmful applications should be explicitly prohibited. Examples include generating misleading or harmful content and developing surveillance systems that infringe on individual privacy rights.
  • Impact Assessment: Users should conduct an impact assessment for their ML programs using Text-to-ML, outlining the potential societal and ethical implications. For example, a project aimed at predicting criminal behavior would need to consider the risks of reinforcing biases and stigmatizing individuals or communities.
Transparency and Accountability in the Automation of Machine Learning
Automating the synthesis of ML programs streamlines the creation of ML models, but it can also obscure the underlying decision-making process. This lack of transparency is particularly concerning in critical fields such as healthcare, finance, criminal justice, job hiring, and college admissions. For instance, in finance, transparency is necessary to ensure fair lending practices and avoid biases in credit scoring. In college admissions, understanding the rationale behind acceptance or rejection decisions is vital for maintaining fairness and avoiding biases that could disadvantage certain groups of applicants. We outline our measures and guidelines to increase transparency and accountability for Text-to-ML as follows:
  • Logging the Automatic Process: The implementation of Text-to-ML in our study features a comprehensive logger that records the entire process of synthesizing the ML program. This includes decisions regarding the search space construction (Appendix C.2), rationales behind the suggestions of candidate solutions (Figure 9), successfully generated modules, and the performance of each module combination being evaluated. We strongly recommend that other implementations of the Text-to-ML framework also incorporate such a logger. This detailed logging information helps users trace outcomes back to specific decisions and increases accountability.
  • Tracing the Prompts and Responses of LLMs: Text-to-ML involves a set of dynamically constructed prompts (Appendix E) in intermediate steps through the instruction constructors ($\{\Phi^{1},\Phi^{2},\ldots,\Phi^{J}\}$ in Algorithm 1). The implementation of Text-to-ML in our study transparently displays the input and output sequences for each inquiry involving LLMs. This history of prompts and responses enables users to understand the decisions made by the LLMs in each step of the process, enhancing transparency and traceability.
  • Interpretable Models: We advocate for the use of interpretable models and techniques that allow users to understand how specific outcomes are derived. This includes the incorporation of explainable AI (XAI) methods (Hoffman et al., 2018) within the Text-to-ML framework to provide insights into the decision-making process.
@@ A.1 Large Language Models
Iterative methods repeatedly generate the output sequences in order to improve the output sequence based on a sequence evaluation procedure. Let the initial input sequence be $X_{0}$, the sequence generation process be $f_{\lambda}$, and the evaluation procedure be $\mathbb{E}$. Then the output sequence in the $i_{th}$ iteration is $Y_{i}=f_{\lambda}(\Phi_{i}(Y_{<i},\mathbb{E}(Y_{<i}),X_{0}))$, where $\Phi_{i}$ represents a mechanism that assembles the input sequence for the iteration by integrating the output sequences, their evaluations, and the initial input sequence from preceding iterations. Recent works along this line include Yao et al. (2022); Madaan et al. (2023); Cobbe et al. (2021); Shinn et al. (2023); Chen et al. (2023b). In a reinforcement learning context, each output sequence represents an action (Yao et al., 2022; Shinn et al., 2023). The implementation of our method incorporates the self-reflection mechanism in Shinn et al. (2023). More specifically, $\Phi_{i}$ in our implementation assembles input sequence from code execution results and an explanation of the results.
Prompting methods improves the output sequence $Y$ by transforming the input sequence $X$. Let $\psi$ represents the prompt engineering process, then $p(Y|X)=\prod_{n=1}^{N}p_{\lambda}(y_{n}|y_{<n},\psi(X))$. Notable prompting methods include in-context learning (Wei et al., 2022; Zhou et al., 2022a; Wang et al., 2022; Min et al., 2022; Yao et al., 2023; Li et al., 2023) and zero-shot prompting (Kojima et al., 2022). Notably, although some prompting methods break the problem into intermediate steps, such as Wei et al. (2022) and Li et al. (2023), these methods perform the steps in one sequence generation process, differing from our definition of compositional reasoning (see Section 2.1).
@@ A.2 Automated Machine learning
Table 4: Comparison with existing methods in terms of qualitative features. (1) Code generation: whether the method automatically generates programs (2) Deep learning: whether the method automatically generates (part of) programs for deep learning tasks. (3) Data preparation: whether the method automatically chooses the strategy for data preparation (4) Modeling: whether the method automatically chooses the strategy for modeling (5) Optimization: whether the method optimizes the programs (components). * Although the method can automatically choose the strategy for modeling, the method does not automatically generate the code for modeling. ** The method uses the LLM as a pseudo-optimization model instead of a dedicated optimization algorithm (e.g., evolutionary computation and reinforcement learning.)
@@ C.1 Unit Testing Details
In this context, distinguishing between the input/output in the ML task and for the modules is crucial. For instance, the input to the data preparation module consists of raw data, while its output comprises the processed input data (features) and output data (targets) in the ML task. The unit tests described in Algorithm 3 focus on the ML task’s input and output data by scrutinizing the output produced by each module when provided with its respective input. For all modules except data preparation, synthetic input/output data are used for testing in the ML task.
@@
In our experiments, we utilize the Python libraries Pytorch, PyTorch Lightning, and Transformers for deep learning tasks and Scikit-learn, XGBoost, CatBoost, and LightGBM for tasks involving tabular data. For deep learning, the specifications in the plan include (1) the tensor types for each tensor. (2) the dimensionality of each tensor. (3) the meaning of each dimension. (4) the size of each dimension. (5) isomorphic dimensions: dimensions that should share the same size over different settings of the data preparation process and modeling configuration. (6) suitable ranges of the size for each dimension. For tabular data tasks, the specifications are limited to the dimensionality of arrays for each of input and output in the ML task.
@@ C.3 Search Strategy
In our study, we employ two search strategies: random search and BOHB (Falkner et al., 2018). Random search is a powerful baseline in hyperparameter optimization due to its wide applicability and high search efficiency for tasks with low effective dimensionality (Bergstra & Bengio, 2012). BOHB combines Bayesian Optimization (BO) and Hyperband (HB) (Li et al., 2018) to optimize hyperparameters efficiently. BOHB inherits HB’s robustness and flexibility, and BO’s ability to ensure strong performance. BOHB starts by rapidly identifying promising configurations like HB and refines them using BO to achieve optimal solutions faster. It’s designed to be simple, computationally efficient, and adaptable to various tasks. For BOHB, we use the default hyperparameters in Falkner et al. (2018) and set 30 epochs as the maximum budget for deep learning tasks.
@@ C.4 Module Generation
In our study, for each ML task, we generate three modules: data preparation, modeling, and post processing. Each module is generated in a step (${j}$) of Algorithm 1. First, the choice specifications (e.g., choice of feature engineering method or choice of models) from Search Space Generation are given as input sequence $X^{j}$. Subsequently, the LLM $f_{\lambda}$ generates the module $Y^{j}$. Then the generated module $Y^{j}$ is assessed with the contextual modular evaluation $\mathbb{E}^{j}$, which is the unit test for the module generated from Algorithm 3. After the tests, Module Generation employs the self-reflection strategy from Shinn et al. (2023) to improve the generation efficiency. The input sequence assembler $\Phi^{j}$ instructs the LLM $f_{\lambda}$ to reflect on the feedback from the tests and then assemble previously generated codes, the tests feedback, and the reflection into the following form for the next iteration:
@@ C.5 ZC proxies details
naswot (Mellor et al., 2021) leverages the overlap of activations between data points in untrained networks to derive a measure indicative of the network’s future trained performance. This approach allows for rapid and efficient searching of powerful networks without the extensive computational costs typically associated with training each candidate network during the search process. Originally designed for network pruning, synflow identifies sparse trainable subnetworks within untrained neural networks. The key principle of synflow is to maintain a balance of synaptic strengths through the network, thus preserving the network’s ability to be trained post-pruning.
All ZC proxies, except for params, require data to be passed to the models for evaluation. To investigate the 3 ZC proxies’ applicability with synthetic data, we first generate the program for producing synthetic data, then sample 50 versions of synthetic data based on different random seeds with a fixed shape of the data. We observe that performance variations reported by each proxy are minimal. Notably, synflow exhibits the most significant variation, with a Mean Relative Deviation (MRD) of $0.76\%$, indicating the suitability of the ZC proxies for synthetic data.
@@ C.7 Hyperparameters for finetuning
For deep learning, our research focuses on an unusual search space with pretrained models at the architectural level. Öztürk et al. (2022) also explored a similar search space. Consequently, we utilize a comparable set of hyperparameters for fine-tuning the models (Table 5). To tailor this approach to our specific requirements, we have reduced the minimum batch size from 16 to 2 to accommodate the varying GPU memory demands across different tasks.
@@ Appendix E Prompts
Next, we present the prompts for the intermediate steps in Text-to-ML . For Figures 23 to 29, texts highlighted in yellow represent prompts based on the common instruction prompt (Figure 10), prompts based on the user input (Figures 11 to 22), or prompts dynamically assembled from responses in previous steps via the instruction constructors ($\{\Phi^{1},\Phi^{2},\ldots,\Phi^{J}\}$ in Algorithm 1. Texts highlighted in green represent prompts that depend on the specific implementation of the pre-written modules (e.g., loss functions, optimizers, training procedure, and testing procedure) and are subject to the constraints in the progenitor testing protocols (Section 4 and Appendix C.1). Texts in red represent prompts configured in the background settings.
Figure 23: Prompt for task analysis
Figure 24: Prompt for task classification
Figure 25: Prompt for data preparation module planning
Figure 26: Prompt for modeling module planning
Figure 27: Prompt for data preparation module code generation
Figure 28: Prompt for modeling module code generation
Figure 29: Prompt for post processing module code generation
@@ Appendix F Additional results
Table 8: Mean iterations to correct errors. (1) tests-only: feedback is provided solely through unit tests (Figure 8 and Algorithm 3). (2) reflection: feedback is provided through both unit tests and the self-reflection mechanism.
Modules GPT-4 GPT-3.5 PaLM 2
tests-only reflection tests-only reflection tests-only reflection
Data preparation 4.1 $\pm$ 1.8 4.2 $\pm$ 1.9 5.6 $\pm$ 2.1 6.0 $\pm$ 1.7 5.8 $\pm$ 2.4 6.1 $\pm$ 2.2
Modeling 2.7 $\pm$ 1.1 2.4 $\pm$ 1.1 2.9 $\pm$ 1.4 3.3 $\pm$ 1.9 5.4 $\pm$ 2.2 5.4 $\pm$ 3.6
Post processing 3.0 $\pm$ 1.0 2.8 $\pm$ 1.4 3.1 $\pm$ 1.9 2.7 $\pm$ 1.2 4.5 $\pm$ 1.5 3.9 $\pm$ 1.2